OpenStudy (anonymous):

Please Help!!!! The apothem of a regular hexagon that has a 36 inch perimeter is _______? I am having a hard time with this!

OpenStudy (lukebluefive):

Does your book give you an equation? I'm having trouble finding one that doesn't involve Trig.

OpenStudy (anonymous):

I will have to find it ! it will take awhile

OpenStudy (anonymous):

@LukeBlueFive i am not seeing any formula for this

OpenStudy (phi):

First, what is the length of one side of your hexagon ?

OpenStudy (anonymous):

@phi would you take the square root of 36 which = 6

OpenStudy (phi):

nice coincidence. The perimeter is the sum of all the sides. there are 6 sides, and the length of them add up to 36. so each has length 36/6 = 6 (regular hexagon means each side has the same length)

OpenStudy (phi):

if the perimeter was 42 inches notice that you do not do sqr(42). you do 42/6 = 7 meaning each side is 7 inches and 7+7+7+7+7+7=42

OpenStudy (phi):



OpenStudy (phi):

any idea what the central angle is ? there are 6 of them, and they add up to 360º (all the way around)

OpenStudy (anonymous):

360/6 =60 is this right

OpenStudy (phi):

yes. Now let's concentrate on one of the 6 triangles